OpenStudy (czarluc):

a group of four students decide to hire a tutor for a review session prior to an examination, and the tutors fee is to be shared by each student. If two additional students join the group, the cost to each student is reduced by 60 Php. How much is the tutors fee?

OpenStudy (anonymous):

You can setup your problem like this \[\frac{ x }{ 4 } - \frac{ x }{ 6 } = 60\] then you just have to solve for x

OpenStudy (anonymous):

from there you can apply the LCM to the left side and solve for x

OpenStudy (czarluc):

why did you equate it to 60?

OpenStudy (anonymous):

you said by 60 php

OpenStudy (anonymous):

I assumed that was your form of dollars

OpenStudy (czarluc):

like how did it become like that?

OpenStudy (czarluc):

yes

OpenStudy (anonymous):

ok. the tutors fee is x and you have 4 students who will split the fee so that gives you x/4 then they say that if you get 2 more students i.e 6 they will split the fee which makes x/4-x/6 now they give you the amount as 60php so \[\frac{ x }{ 4 } - \frac{ x }{ 6 } = 60\]

OpenStudy (anonymous):

now if you use the LCM you'll get \[\frac{ 3 }{ 3 } * \frac{ x }{ 4 } - \frac{ 2 }{ 2 } *\frac{ x }{ 6 } = 60\] then you just combine and solve for x

OpenStudy (anonymous):

I hope that makes sense..

OpenStudy (czarluc):

thanks

OpenStudy (anonymous):

:)